Fame is 10 yrs. older than EJ. Five years ago, EJ was two years more than one fourth as old as Fame. What are their present ages?
Answer by stanbon(75887)

Fame is 10 yrs. older than EJ. Five years ago, EJ was two years more than one fourth as old as Fame. What are their present ages?
--------------
Equations:
f = e + 10
(e-5) = (1/4)(f-5)+2
----
Substitute for "f" and solve for "e":
e-5 = (1/4)(e+10-5)+2
e-5 = (1/4)(e+5)+2
Multiply thru by 4 to get:
4e-20 = e+5+8
3e = 33
e = 11 yrs old (EJ's age now)
---
Since f = e+10, f = 21 yrs old (Fame's age now)
=========================
Cheers,
Stan H.
=======================